# OORD: The Oxford Offroad Radar Dataset

### v1

@@ VII Results
As can be seen in Tab. III, considering that all of each of the routes is at minimum in length (seeTab. II), all radar-specific methods perform very well, with in excess of for even Recall@1 i.e. when relying on only a single nearest neighbour in embedding space look up.
@@ VIII Conclusion
We have presented a novel radar dataset, carefully gathered under unique and challenging conditions. This dataset has been crafted to catalyse advancements in the emerging field of radar place recognition. We explored the use of our dataset in this task over a series of comprehensive experiments and evaluations, carried out across various open-source radar place recognition systems. This was not only to showcase the dataset but also to establish a robust platform for future research in this evolving area.

### v2

@@ VII Results
VII-A Baseline comparison
As can be seen in Tab. III, considering that all of each of the routes is at minimum in length (seeTab. II), all radar-specific methods perform very well, with in excess of for even Recall@1 i.e. when relying on only a single nearest neighbour in embedding space look up. Tab. III also shows results when using and nearest neighbours. Additionally, Fig. 9 shows results when using all of to nearest neighbours, as “recall curves” rather than tabulated.
@@ VII-B Failure cases
VII-C Comparison against visual place recognition
Finally, Fig. 10 shows a comparison between visual place recognition and radar place recognition on this dataset. We see that radar has better performance in this task on this dataset across both experimental settings. For example, we have almost recall at any number of database candidates () – see the orange curve – while a vision-based approach has less than Recall@1 and this rises to only about (red curve). This was an experiment on LABEL:DB1 versus LABEL:DB2, with little appearance variation (see Fig. 7, top left) The next experiment (blue and green) is on LABEL:DM1 versus LABEL:DM2, which has a lot of appearance variation (see Fig. 7 where LABEL:DM2 is under very poor lighting – with driving taking place at night, in the dark). Here, radar does only very slightly worse – with over Recall@1 (Fig. 10 leftmost blue dot and Tab. III third row, fourth column). Performance with the camera, however, is drastically reduced – e.g. dropping to approximately Recall@1. This is an important result as it confirms that radar is somewhat immune to extreme appearance variation and is a useful counterpart (for localisation in this case) to the comparison to [32] where radar was shown to be quite robust to appearance change in urban environments in the task of vehicle detection and tracking.
@@ VIII Conclusion
We have presented a novel radar dataset, carefully gathered under unique and challenging conditions. This dataset has been crafted to catalyse advancements in the emerging field of radar place recognition. We explored the use of our dataset in this task over a series of comprehensive experiments and evaluations, carried out across various open-source radar place recognition systems. This was not only to showcase the dataset but also to establish a robust platform for future research in this evolving area. We found that the best performing radar place recognition methods in terms of successful localisation queries over thousands of trials over pairs of sequences from distinct routes were the recently available open-source systems OpenRadVLAD, RaPlace, and ResNet18-NetVLAD the last of which is released alongside this dataset. We also proved the usefulness of the dataset in confirming appearance-change robustness for radar in comparison to visual sensors.